# Enabling decision support over confidential data

### v1

@@ IV-B1 System initialization
Table III: Process participants’ and CCU types and attributes
@@
In this step, the CA provides evidence of the attributes held by actors using a certificate. Table III lists the types and attributes of the Data Providers involved in Sect. III and of the CCU component, too. For example, the CA attests that the medical hub who is both a Data Provider and a Decider, holds attributes specifying an IdentificationNumber, and that it is a MedicalHub operating in Italy. This certificate contains the Elliptic Curve Digital Signature Algorithm (ECDSA) [39] public key of the actor and their attributes. Issuing the certificate, the CA attests to this information for each user. In our approach, we implemented a CA that simulates the issuance of X.509 certificates, signing them with its ECDSA private key. Listing 1 shows an example of such a certificate. Each participant must check the validity of the counterpart’s certificate with the CA’s ECDSA public key before any interaction (as per Tab. II). Furthermore, the certificate has a unique identification number, a time validity, and the user’s ECDSA public key. Notably, the CCU must also obtain a valid certificate from the CA to let the users check whether they are interacting with the correct computation unit (meeting Tab. II).
@@ 10 IDCheck=True
Before making any request, a process participant like the Data Provider, the Decider, or even other CCUs (henceforth, a client) performs the remote attestation to establish trust with the CCU deployed inside a TEE (as per Tab. II). This phase follows the principles defined in the RATS RFC standard [40], the foundation for several TEE attestation schemes, such as Intel EPID and AMD SEV-SNP. Remote attestation serves a dual objective: on one hand, provide participants with strong evidence that the data they are transmitting is sent to a Trusted Application running within a TEE; on the other hand, verify the identity and integrity of the DeSObj as an authentic CCU software entity, recognized as such by the CA (see the CERT step) within the context of the running process (in this case, the scenario described in Sect. III). We slightly modify the remote attestation, to meet all the features described. To perform it, the client (i) retrieves and checks the CCU’s certificate released by the CA to verify that the CCU is the intended one and not a malicious attacker. Then, (ii) the client retrieves and validates the TLS certificate issued by the CCU to establish a secure connection with it, and finally (iii) gets (via attested TLS channel) and verifies the attestation report of the CCU. The attestation report contains the so-called measurement of the Trusted App, which is defined as the hash value of the combination of its source code and data. This report is signed using the attestation private key associated with the CCU’s hardware. To verify it, the client (a) authenticates the CCU using the public key and decrypts the report, (b) compares the measurement found within the decrypted report against a predefined reference value associated with the source code of the CCU (if the decrypted measurement matches the predefined value, the CCU gains trust from the client), (c) verifies that the report data matches the server’s Transport Layer Security (TLS) certificate.
@@
Relying on a single CCU introduces significant vulnerabilities that could compromise Tabs. II and II. If this component becomes unavailable due to failure, downtime, or malicious attacks, the entire execution process could halt indefinitely, potentially making it impossible to recover. Additionally, if that CCU exclusively holds the cryptographic keys, all associated data could be permanently lost. We recall that the Request Manager is not within a TEE, so information exchanges with the system should be encrypted with a known key Sect. IV-A. To deal with these issues in SPARTA, the CCU creates and shares with other CCUs a seed to ensure that they all generate the necessary cryptographic keys deterministically (meeting Tab. II). To explain how we achieve this, let us consider a scenario with two CCUs: CCU1 and CCU2. Let us assume that CCU1 generates the seed and sends it to CCU2. CCU1 does so as follows. (1) It performs the remote attestation of CCU2. If the remote attestation (RA ) succeeds, (2) it generates a secure random string (the seed) and seals it with a sealing key known only to itself (see Sect. II); then, it sends the seed to CCU2 via the secure channel established during RA . As CCU2 receives the secret seed from CCU1, it performs the first two operations as well, namely it performs the remote attestation of CCU1, and, if the latter succeeds, it seals the seed with a sealing key known only to itself. At that point, the exchanged seed transmitted by CCU1 can be shared with an arbitrary number of other CCUs. We will describe how the seed is used to generate encryption keys next, and in Sect. IV-B2.
@@
After the seed exchange, a CCU can generate a certificate for establishing secure communication channels to and from the Request Manager via TLS over TCP, so that confidential data is not leaked (as per Tab. II). More specifically, we slightly modify the TLS certificate, allowing it to contain, as an extra parameter, an Elliptic-Curve Diffie-Hellman (ECDH) [41] public key. The CCU loads the seed created in the SDEX step and uses it to generate a deterministic ECDH private/public key pair (using the X25519 curve [42]). Notice that with more than one CCU, since all of them share the same seed (see the SDEX step), they generate the same ECDH key pair. This ECDH key is used by the client to generate a shared secret before transmission to the Request Manager (see Sect. IV-B2). Since the CCU signs the TLS certificate with its ECDSA private key and the corresponding ECDSA public key is present inside the certificate, every participant can verify the signature with the CCU’s ECDSA public key. The ECDSA public key is certified by the CA (see the CERT step), so a participant can trust both the certificate and the CCU. At the end of this step, the CCU is ready to receive requests and provide responses through the Request Manager.
@@
After the certification phase, the Policymaker declares the decision functions the system should realize in a decision model, and their access rules in a policy descriptor. In order to meet Tab. II, the Policymaker should be allowed to express those with standard, well-established, and customizable languages. Here we resort to DMN with FEEL, and ALFA, respectively (see Sect. II). An explanation of DMN, FEEL, and ALFA scripting languages transcends the scope of this paper. It suffices to explain their use here by example.
Table IV: Decision model for the patient prioritization of Sect. III
@@
Table IV shows an example of a DMN indicating how a decision is made for the prioritization of vaccine requests by patients. All input columns refer to data attributes in the structure retaining patient information. Among them, the meanAge and sumAge columns are aggregation functions. For the sake of readability, we separately show their encoding in FEEL in Tab. V. Considering the meanAge aggregation, e.g., the FEEL formula filters the data (based, among others, on the patients’ family medical history, age, and previous vaccinations), and then computes the mean age of the remaining entries. The DeFObj builder (see Sect. IV-A) transforms such a table into executable code.
@@
Table V: FEEL formulas the functions in List. 3
@@
Input: descriptorFile containing: funcName, the string representing the name of the function; accPol, the function’s access policy; datStNam, the indexing name of the input data structure; aggrFuncs, a (possibly empty) list of aggregation functions on the input data.
@@
Algorithm 1 Code generator for DeSObj instances
@@
Notice that this script does not contain direct calls to the decision functions, or data retrieval. It delegates the actual execution to the DeLib instead. The code inside the DeLib is parametric and independent of the use case scenario. Only the DeSObj script is tailored to the specific decision process. Finally, the DeTApp builder deploys the above software objects within the TEE enclave as a DeTApp.
@@
IV-B2 Execution
The lilac box encloses the components involved in the execution phase of SPARTA Fig. 1. It encompasses a writing phase, and a decision making phase. In the writing stage (WS) , data is provided for decisions, meeting Tabs. II and II. It is comprised of two steps, respectively corresponding to the data provision and data usage stages in our motivating scenario (Sect. III).
@@ IV-B2 Execution
This step basically establishes a secure asynchronous communication channel over an insecure channel, i.e., the Request Manager’s queue, between the client and the CCU. It pertains to the Data Provider when sending the input data to the CCU through the Request Manager. As specified in Sect. IV-A, the Request Manager is not trusted, so data must be encrypted before its transmission; otherwise, another participant or a malicious attacker can sniff data from the queue. To avoid this, the Data Provider, when performing the remote attestation against the CCU, retrieves the CCU’s ECDH public key from the TLS certificate (see the CONF step). Then, they generate an ECDH private/public key pair [41] using the X25519 curve [42] and use their own private ECDH key along the CCU’s ECDH public key to generate a shared secret thanks to the ECDH algorithm. From the shared secret, the Data Provider generates a symmetric key deterministically by using Secure Hash Algorithm (SHA) (in particular SHA-256 [43, 44]) and uses this key to encrypt the data before sending it to the CCU via the Request Manager using AES-GCM [45, 46]. Moreover, the Data Provider signs their own ECDH public key with their ECDSA private key (see the CERT step) and attaches the signed ECDH public key and the ECDSA public key in the payload. After reading the input data from the request queue, the CCU checks the CA certificate of the invoking Data Provider (using the ECDSA public key of the CA), then checks the ECDH signature with the ECDSA public key (of the Data Provider), and finally generates the shared secret thanks to its private ECDH key and the Data Provider’s public ECDH key with the ECDH algorithm. The shared secret will be equal to the one calculated by the Data Provider in the encryption phase. From that secret, it generates the symmetric key, and since this operation is deterministic, this key will also be equal to the one calculated by the Data Provider. At this point, the CCU decrypts the data sent by the Data Provider. The symmetric key exchange mechanism described above guarantees that the Request Manager, even if fully corrupted, cannot learn the key shared between the Data Provider and the CCU, thus ensuring that the Request Manager cannot read any message encrypted using a secure symmetric key encryption mechanism such as AES under the shared symmetric key.
@@
The second writing phase pertains to the CCU and follows the Data Provider’s input data decryption. To securely store data in a public IPFS, the CCU generates a new unique symmetric key from a pre-determined private seed each time a new element must be saved to IPFS. Since these keys are deterministically computed from the CCU’s private seed, the CCU can easily recompute the very same key each time a particular piece of data must be decrypted for computing the final decisions. In practice, to compute this derived symmetric key, the CCU generates a random number and calculates the SHA-3 hash [43] of the seed shared during the SDEX step, and the random number. Finally, it uses this symmetric key to encrypt the data with AES-GCM (meeting Tab. II,Tab. II). The encrypted data is put on IPFS. The random number must be freshly generated for each record and must be the same as the one used in the decryption phase described next. The fact that each random number for each entry is written on IPFS in clear does not impact security since the SHA-3 operation on each random number and the seed produces a new symmetric key which is totally unrelated from the CCU’s private seed and the other keys. To meet Tab. II, the CCU stores all the necessary IPNS names in a smart contract on a public blockchain. It is noteworthy to say that since the CCU encrypts data also using a seed established during the SDEX step, malicious attackers who were not part of this phase cannot impersonate a CCU to access the data, as they cannot steal the seed to reconstruct the decryption key because it is encrypted with a sealing key (see Sect. II)—and they cannot be part of the SDEX procedure without a valid certificate form the CA.
@@
This phase corresponds to the data usage stage described in Sect. III. When a Decider asks for decision support, they invoke the decision function of the DeSObj deployed inside the CCU as per Tab. II (e.g., the PatientPrioritizationWithAggrHandler in List. 4). Recall that every request is handled by the Request Manager. Before sending the request, the participant performs the remote attestation against the CCU (see the note on RA). If the attestation completes successfully, the participant finalizes the request. The payload of the request contains the Decider’s certificate released by the CA, the name of the function they are invoking (in this case, PatientPrioritizationWithAggr), and the IPNS key name (e.g., PatientsWAggrKey). As in List. 4, after parsing the participant request with the DeLib (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:payload), the CCU checks the received CA certificate of the requester invoking the DeLib (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:check-certificate). If it is valid, then the CCU extracts the attributes inside the certificate and checks whether they satisfy the function’s access policy linked to the invoked decision function thanks to the DeLib (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:check-callability). In our example, the access policy is isMedicalHub. If the attributes of the Decider are the ones of List. 1, the access policy of the invoked function is satisfied by the attributes the Decider bears, so they can invoke it. At this stage, the CCU retrieves and decrypts the data with the corresponding function inside the DeLib (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:decrypt-data) and performs the two aggregations functions evaluating the corresponding FEEL formulas invoking the DeLib (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:mean-aggregationList. 4, ln. LABEL:lst:FinalDecisionWithAggregation:sum-aggregation). Finally, it invokes the desired decision function (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:decision, note that the PatientPrioritizationWithAggr decision function corresponds to the List. 2). Finally, the CCU sends the result back to the Request Manager that returns it to the Decider (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:return).
V Formal Analysis

### v2

@@ 4.2.1 System initialization
Table III: Types and attributes of participants and CCU
@@
In this step, the CA provides evidence of the attributes held by actors using a certificate. Table III lists the types and attributes of the Data Providers involved in Sect. 3 and of the CCU component, too. For example, the CA attests that the medical hub who is both a Data Provider and a Decider, holds attributes specifying an IdentificationNumber, and that it is a MedicalHub operating in Italy. This certificate contains (among other metadata) the Elliptic Curve Digital Signature Algorithm (ECDSA) [ECDSA] public key of the actor and their attributes. Issuing the certificate, the CA attests to this information for each user. In our approach, we implemented a CA that simulates the issuance of X.509 certificates, signing them with its ECDSA private key. Listing 1 shows an example of such a certificate. Each participant must check the validity of the counterpart’s certificate with the CA’s ECDSA public key before any interaction (as per Tab. II). Notably, the CCU must also obtain a valid certificate from the CA to let the other actors involved to check whether they are interacting with the correct computation unit (meeting Tab. II). To perform signatures, actors use their ECDSA private key related to the public one inside the certificate attested by the CA. Otherwise, the signature is not valid.
@@ 10 IDCheck=True
Before making any request, a process participant like the Data Provider, the Decider, or even other CCUs (henceforth, a client) performs the remote attestation to establish trust with the CCU deployed inside a TEE (as per Tab. II). This phase follows the principles defined in the RATS RFC standard [birkholz2023rfc], the foundation for several TEE attestation schemes, such as Intel EPID and AMD SEV-SNP. Remote attestation serves a dual objective: on one hand, provide participants with strong evidence that the data they are transmitting is sent to a Trusted Application running within a TEE; on the other hand, verify the identity and integrity of the DeSObj as an authentic CCU software entity, recognized as such by the CA (see the CERT step) within the context of the running process (in this case, the scenario described in Sect. 3). We slightly modify the remote attestation, to meet all the features described. To perform it, the client (i) retrieves and checks the CCU’s certificate issued by the CA using CA’s ECDSA public key, verifying that the CCU is the intended one and not a malicious attacker. Then, (ii) the client retrieves and validates the Transport Layer Security (TLS) certificate issued by the CCU (checking also CCU’s ECDSA public key among other things) to establish a secure connection with it, and finally (iii) gets (via attested TLS channel) and verifies the attestation report of the CCU. This report contains the so-called measurement of the Trusted App, which is defined as the hash value of the combination of its source code and data. This report is signed using the attestation private key associated with the CCU’s hardware. To verify it, the client (a) authenticates the CCU using the public key and decrypts the report, (b) compares the measurement found within the decrypted report against a predefined reference value associated with the source code of the CCU (if the decrypted measurement matches the predefined value, the CCU gains trust from the client), (c) verifies that the report data matches the server’s TLS certificate.
@@
Relying on a single CCU introduces significant vulnerabilities that could compromise Tabs. II and II. If this component becomes unavailable due to failure, downtime, or malicious attacks, the entire execution process could halt indefinitely, potentially making it impossible to recover. Additionally, if that CCU exclusively holds the cryptographic keys, all associated data could be permanently lost. We recall that the Request Manager is not within a TEE, so information exchanges with the system should be encrypted with a known key Sect. 4.1. To deal with these issues in SPARTA, the CCU creates and shares with other CCUs a seed to ensure that they all generate the necessary cryptographic keys deterministically (meeting Tab. II). To explain how we achieve this, let us consider a scenario with two CCUs: CCU1 and CCU2. Let us assume that CCU1 generates the seed and sends it to CCU2. CCU1 does so as follows. (1) It performs the remote attestation of CCU2. If the remote attestation (RA ) succeeds, (2) it generates a secure random string (the seed) and seals it with a sealing key known only to itself (see Sect. 2); then, it sends the seed to CCU2 via the secure channel established during RA . As CCU2 receives the secret seed from CCU1, it performs the first two operations as well, namely it performs the remote attestation of CCU1, and, if the latter succeeds, it seals the seed with a sealing key known only to itself. At that point, the exchanged seed transmitted by CCU1 can be shared with an arbitrary number of other CCUs. We will describe how the seed is used to generate encryption keys next, and in Sect. 4.2.2.
@@
After the seed exchange, a CCU can generate a certificate for establishing secure communication channels to and from the Request Manager via TLS over TCP, so that confidential data is not leaked (as per Tab. II). More specifically, we slightly modify the TLS certificate, allowing it to contain, as an extra parameter, an Elliptic-Curve Diffie-Hellman (ECDH) [ECDH] public key. The CCU loads the seed created in the SDEX step and uses it to generate a deterministic ECDH private/public key pair (using the X25519 curve [Curve25519]). Notice that with more than one CCU, since all of them share the same seed (see the SDEX step), they generate the same ECDH key pair. This ECDH key is used by the client to generate an exchanged secret before transmission to the Request Manager (see Sect. 4.2.2). Since the CCU signs the TLS certificate with its ECDSA private key and the corresponding ECDSA public key is present inside the certificate, every participant can verify the signature with the CCU’s ECDSA public key. The ECDSA public key is certified by the CA (see the CERT step), so a participant can trust both the certificate and the CCU. At the end of this step, the CCU is ready to receive requests and provide responses through the Request Manager.
@@
After the certification phase, the Policymaker declares the decision functions the system should realize in a decision model, and their access rules in a policy descriptor. In order to meet Tab. II, the Policymaker should be allowed to express those with standard, well-established, and customizable languages. We resort to DMN with FEEL, and ALFA, respectively (see Sect. 2). An explanation of DMN, FEEL, and ALFA scripting languages transcends the scope of this paper. It suffices to explain their use by example.
Table IV: Decision model for the patient prioritization of Sect. 3
@@
Table IV shows an example of a DMN indicating how a decision is made for the prioritization of vaccine requests by patients. All input columns refer to data attributes in the structure retaining patient information. Among them, the meanAge and sumAge columns are aggregation functions. For the sake of readability, we separately show their encoding in FEEL in Tab. V. Considering the meanAge aggregation, e.g., the FEEL formula filters the data (based, among others, on the patients’ family medical history, age, and previous vaccinations), and then computes the mean age of the remaining entries. The DeFObj builder (see Sect. 4.1) transforms such a table into executable code.
@@
Table V: FEEL formulas of the functions in List. 3
@@
Input: A descriptorFile containing: funcName, the string representing the name of the function; accPol, the function’s access policy; datStNam, the indexing name of the input data structure; aggrFuncs, a (possibly empty) list of aggregation functions on the input data.
@@
Algorithm 1 Code generator for DeSObj
@@
Notice that this script does not contain direct calls to the decision functions, or data retrieval. It delegates the actual execution to the DeLib instead. The code inside the DeLib is parametric and independent of the use case scenario. Only the DeSObj script is tailored to the specific decision process. Finally, the DeTApp builder deploys the software objects within the TEE enclave as a DeTApp.
@@
4.2.2 Execution
The lilac box encloses the components involved in the execution phase of SPARTA Fig. 1. It encompasses a writing phase, and a decision making phase. In the writing stage (WS) , data is provided for decisions, meeting Tabs. II and II. It is comprised of two steps, respectively corresponding to the data provision and data usage stages in our motivating scenario (Sect. 3).
@@ 4.2.2 Execution
This step basically establishes a secure asynchronous communication channel over an insecure channel, i.e., the Request Manager’s queue, between the client and the CCU. It pertains to the Data Provider when sending the input data to the CCU through the Request Manager. As specified in Sect. 4.1, the Request Manager is not trusted, so data must be encrypted before its transmission; otherwise, another participant or a malicious attacker can sniff data from the queue. To avoid this, the Data Provider, when performing the remote attestation against the CCU, retrieves the CCU’s ECDH public key from the TLS certificate (see the CONF step). Then, the Data Provider generates an ECDH private/public key pair [ECDH] using the X25519 curve [Curve25519] and use their own private ECDH key along the CCU’s ECDH public key to generate an exchanged secret thanks to the ECDH algorithm. From this exchanged secret, the Data Provider generates a symmetric key deterministically by using Secure Hash Algorithm (SHA) (in particular SHA-256 [SHA, SHA2]) and uses this key to encrypt the data before sending it to the CCU via the Request Manager using AES-GCM [AES-GCM1, AES-GCM2]. Moreover, the Data Provider signs their own ECDH public key with their ECDSA private key (see the CERT step) and attaches the signed ECDH public key and the ECDSA public key in the payload. After reading the input data from the request queue, the CCU checks the CA certificate of the invoking Data Provider (using CA’s ECDSA public key), then checks the ECDH signature with Data Provider’s ECDSA public key, and finally generates the exchanged secret thanks to its private ECDH key and the Data Provider’s public ECDH key with the ECDH algorithm. The exchanged secret will be equal to the one calculated by the Data Provider in the encryption phase. From that secret, the CCU generates the symmetric key, and since this operation is deterministic, this key will also be equal to the one calculated by the Data Provider. At this point, the CCU decrypts the data sent by the Data Provider. The symmetric key exchange mechanism described above guarantees that the Request Manager, even if fully corrupted, cannot learn the key shared between the Data Provider and the CCU, thereby ensuring that it cannot decrypt any message encrypted with a secure symmetric scheme under that shared key.
@@
The second writing phase pertains to the CCU and follows the Data Provider’s input data decryption. To securely store data in a public IPFS, the CCU generates a new unique derived symmetric key from a pre-determined private seed each time a new element must be saved to IPFS. Since these keys are deterministically computed from the CCU’s private seed, the CCU can easily recompute the very same key each time a particular piece of data must be decrypted for computing the final decisions. In practice, to compute this derived symmetric key, the CCU generates a random number and calculates the SHA-3 hash [SHA] of the seed shared during the SDEX step, and the random number. Finally, it generates a symmetric key starting from the SHA-3 result, and uses this symmetric key to encrypt the data with AES-GCM (meeting Tab. II,Tab. II). The encrypted data is put on IPFS. The random number must be freshly generated for each record and must be the same as the one used in the decryption phase described next. The fact that each random number for each entry is written on IPFS in clear does not impact security since the SHA-3 operation on each random number and the seed produces a new symmetric key which is totally unrelated from the CCU’s private seed and the other keys. To meet Tab. II, the CCU stores all the necessary IPNS names in a smart contract on a public blockchain. It is noteworthy to say that since the CCU encrypts data also using a seed established during the SDEX step, malicious attackers who were not part of this phase cannot impersonate a CCU to access the data, as they cannot steal the seed to reconstruct the decryption key because it is encrypted with a sealing key (see Sect. 2)—and they cannot be part of the SDEX procedure without a valid certificate form the CA. Additionally, if a CCU is unavailable, another CCU that took part in the SDEX step can take its place, since this derived symmetric key is also generated from the shared seed.
@@
This phase corresponds to the data usage stage described in Sect. 3. When a Decider asks for decision support, they invoke the decision function of the DeSObj deployed inside the CCU as per Tab. II (e.g., the PatientPrioritizationWithAggrHandler in List. 4). Recall that every request is handled by the Request Manager. Before sending the request, the participant performs the remote attestation against the CCU (see the note on RA). If the attestation completes successfully, the participant finalizes the request. The payload of the request contains the Decider’s certificate released by the CA, the name of the function they are invoking (in this case, PatientPrioritizationWithAggr), and the IPNS key name (e.g., PatientsWAggrKey). As in List. 4, after parsing the participant request with the DeLib (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:payload), the CCU checks the received CA certificate of the requester (using CA’s ECDSA public key) invoking the DeLib (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:check-certificate). If it is valid, then the CCU extracts the attributes inside the certificate and checks whether they satisfy the function’s access policy linked to the invoked decision function thanks to the DeLib (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:check-callability). In our example, the access policy is isMedicalHub. If the attributes of the Decider are the ones of List. 1, the access policy of the invoked function is satisfied by the attributes the Decider bears, so they can invoke it. At this stage, the CCU retrieves and decrypts the data (reconstructing the derived symmetric key) with the corresponding function inside the DeLib (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:decrypt-data) and performs the two aggregations functions evaluating the corresponding FEEL formulas invoking the DeLib (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:mean-aggregation, List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:sum-aggregation). Finally, it invokes the desired decision function (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:decision, note that the PatientPrioritizationWithAggr decision function corresponds to the List. 2). Finally, the CCU sends the result back to the Request Manager that returns it to the Decider (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:return).
Table VI: Keys of CA, Data Provider, and CCU
Subject Key Function
CA ECDSA key pair Uses private key to sign certificates (Listing 1) in CERT
Others use public to verify certificate’s signature in RA, WS-i, DECM
Data Provider ECDSA key pair Uses private to sign ECDH public key in WS-i
CCU uses public to verify ECDH public key signature in WS-i
ECDH key pair Uses private to generate exchanged secret (with CCU’s public) in WS-i
CCU uses public to generate the same exchanged secret in WS-i
Symmetric key Generated from exchanged secret to encrypt data with AES-GCM in WS-i
CCU ECDSA key pair Uses private to sign TLS certificate in CONF
Others use public to verify TLS certificate signature in RA
ECDH key pair Key pair generated starting from the seed (SDEX) in CONF
Uses private to generate exchanged secret (with Data Provider’s public) in WS-i
Data Provider uses public to generate the same exchanged secret in WS-i
Symmetric key Generated from exchanged secret to decrypt data with AES-GCM in WS-i
Sealing key Uses it to encrypt seed in SDEX and decrypt it in WS-ii
Derived Generated from seed (SDEX) and a random number
Symmetric Key to encrypt data with AES-GCM in WS-ii and decrypt it in DECM
5 Formal Analysis